if lucy was three times as old as she was four years ago she will be fourteen more than her current age. how old is lucy
now

Answer by ankor@dixie-net.com(22740)

if Lucy was three times as old as she was four years ago she will be fourteen more than her current age. how old is Lucy
now
let a = Lucy's present age
then
(a-4) = Her age 4 yrs ago
:
Write an equation for just what it says
3(a-4) = a + 14
3a - 12 = a + 14
3a - a = 14 + 12
2a = 26
a = 26/2
a = 13 yrs is her present age
:
You can confirm this by replacing a with 13 in the original equation